OpenStudy (anonymous):

1.) 6 sqrt.5 ( [srt.11] + [sqrt.5] ) 2.) (3 sqrt.3 +3 )(6 sqrt. 3 + 2)

OpenStudy (anonymous):

1.) \[6\sqrt{5} ( \sqrt{11} + \sqrt{5})\]

OpenStudy (amistre64):

what do you come up with when you distribute it thru?

OpenStudy (anonymous):

2.) \[(3\sqrt{3} + 3) (6\sqrt{3} + 2)\]

OpenStudy (anonymous):

what do you mean ?

OpenStudy (amistre64):

I mean that you need to know what the distributive property is so you can solve these for yourself ;)

OpenStudy (amistre64):

if you want to learn, I can teach you :)

OpenStudy (amistre64):

and posting 2 questions at a time tends to mean you just want someone else to do your homework for you.....

OpenStudy (anonymous):

\[2.) 6 + 18\sqrt{3}^{2} + 6\sqrt{3} \]

OpenStudy (anonymous):

i think thats the answer

OpenStudy (anonymous):

and no its not homework its an exam for my course

OpenStudy (amistre64):

its a good start.... but lets check it out...

OpenStudy (anonymous):

i dont think i can simplify it more ?

OpenStudy (amistre64):

(3 sqrt.3 +3 )(6 sqrt. 3 + 2) 3sqrt(3) (6sqrt(3)) = 18(3) 3(2)sqrt(3) = 6sqrt(3) 3(6sqrt(3)) = 18sqrt(3) 3(2)=6

OpenStudy (anonymous):

so im right ? :)

OpenStudy (amistre64):

54+6 +24sqrt(3) 60 + 24sqrt(3) if I did it right :)

OpenStudy (amistre64):

you had 3 terms good, but were missing that forth

OpenStudy (amistre64):

12(5 + 2sqrt(3)) is another way to write it

OpenStudy (anonymous):

ohhh true ! and for the 1st problem i got \[6\sqrt{11} + 5\]

OpenStudy (anonymous):

im not sure if its right ?

OpenStudy (amistre64):

I got: 6sqrt(55) + 30

OpenStudy (amistre64):

\[6\sqrt{5*11} + 6\sqrt{5*5}\]

OpenStudy (anonymous):

ahhh true , i thought i was just suppose to cancel it out

OpenStudy (amistre64):

\[6\sqrt{55} + 6(5) = 6\sqrt{55}+30\]

OpenStudy (anonymous):

\[(2x \sqrt{12x}) (3\sqrt{4x})\]

OpenStudy (anonymous):

can you help me with that problem aswell ?

OpenStudy (anonymous):

i got \[24x ^{2}\sqrt{3}\]

OpenStudy (amistre64):

multiplication doesnt care in what order you do it...\[2x(3) *\sqrt{12x}*\sqrt{4x} = 6x \sqrt{48x^2}\]

OpenStudy (anonymous):

cant i simplify that further ?

OpenStudy (amistre64):

6x 4x sqrt(3) 24x^2 sqrt(3)

OpenStudy (anonymous):

same as my answer ?

OpenStudy (amistre64):

you were right :)

OpenStudy (anonymous):

thank youuu :) i had to double check ! :D how about this one : \[\sqrt{2x ^{2}} \sqrt{10x ^{6}}\]

OpenStudy (anonymous):

i got \[2x ^{4} \sqrt{5}\]

OpenStudy (radar):

You seem to have the hang of it!

OpenStudy (anonymous):

so its right ? :)

OpenStudy (radar):

It is the same as what I got. So I think it is right.

OpenStudy (anonymous):

and its the same steps for addition ? for example \[4^{3}\sqrt{a} + ^{3}\sqrt{64a}\]

OpenStudy (anonymous):

i got \[16^{3}\sqrt{a}\]

OpenStudy (anonymous):

but im not sureee

OpenStudy (radar):

Is that 4 cuberoot of a + cube root of 64a? If it is\[4\sqrt[3]{a}+\sqrt[3]{64a}=8\sqrt[3]{a}\]

OpenStudy (anonymous):

yes it is , your right . i got confused with it

OpenStudy (radar):

You're doing fine, I've got to run. good luck

OpenStudy (anonymous):

thank youuuu ! :)